OpenStudy (anonymous):

. Heather has a bag with 2 marbles, 8 tennis balls, 3 apples, and 4 books. What is the ratio of books to tennis balls? (8/4) or 8:4 (3/7) or 3:7 (4/8) or 4:8 (2/8) or 2:8

OpenStudy (anonymous):

Help

OpenStudy (texaschic101):

books to tennis balls how many book you got ? how many tennis balls you got ?

OpenStudy (anonymous):

4 books and 8 tennis balls

OpenStudy (texaschic101):

so the ratio is 4:8 or 4/8 you have to pay attention to the wording because if it said tennis balls to books, it would be 8/4 or 8:4

OpenStudy (texaschic101):

understand ?

OpenStudy (anonymous):

would the small number come before the bigger number

OpenStudy (anonymous):

brb

OpenStudy (texaschic101):

depends on the ratio books to tennis balls = 4/8 or 4:8 tennis balls to books = 8/4 or 8:4 so this answer would be 4:8 or 4/8 because you are looking for book to tennis balls

OpenStudy (anonymous):

ok thanks

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

i gave u a medal